# A Comprehensive Survey on Knowledge Distillation

### v1

@@ V-B Multi-view Input
TABLE VII: Summary of Knowledge Distillation in Multi-View Tasks.
@@
KD-MVS[331] (2022) Multi-view Stereo Depth Reconstruction Multi-view Images Multi-view Images Self-supervised MVS distillation for depth reconstruction.
@@
MTS-Net[332] (2022) Multi-view Image Classification Multi-view Images Multi-view Images Multi-view learning and distillation for image classification.
@@
BEV-LGKD[333] (2022) Multi-view BEV 3D Object Detection LiDAR Multi-view RGB Images LiDAR-guided distillation with foreground masks and depth distillation for BEV perception.
@@
BEVDistill[334] (2022) Multi-view 3D Object Detection LiDAR Multi-view Images Cross-modal BEV distillation for unifying image and LiDAR features.
@@
STXD[335] (2023) Multi-view 3D Object Detection LiDAR Multi-view Images Structural and temporal distillation with cross-correlation regularization and response distillation.
@@
PointMCD[313] (2023) 3D Shape Recognition Multi-view Images 3D Point Cloud Multi-view cross-modal distillation for 3D shape recognition.
@@
Acar et al.[336] (2023) Vision-based Robotic Manipulation Multi-view Images Single-camera Distillation from multi-camera to single-camera for robust manipulation.
@@
Zhang et al.[337] (2023) Multi-view BEV Detection Multi-view BEV Images Multi-view BEV Images Structured distillation for efficient BEV detection.
@@
MVKD[338] (2023) Semantic Segmentation Multi-view Images Single-view Images Multi-view distillation with co-tuning and feature/output distillation losses.
@@
MKDT[339] (2023) Human Action Recognition Multi-view Videos Single-view Videos Multi-view distillation with hierarchical vision transformer for incomplete data.
@@
3X3KD[210] (2023) Multi-camera 3D Object Detection LiDAR Multi-view Images Cross-modal distillation for multi-camera 3D object detection.
@@
DistillBEV[209] (2023) 3D Object Detection for Autonomous Driving LiDAR Multi-view BEV Images Distillation from LiDAR to BEV for enhanced 3D perception.
@@
FSD-BEV[340] (2024) 3D Object Detection for Autonomous Driving Self-distillation Self-distillation Foreground self-distillation for single-model distillation.
@@
SimDistill[341] (2024) Multi-modal 3D Object Detection LiDAR-camera Fusion Multi-view Images Distillation for 3D object detection with modality gap compensation.
@@
GMViT[342] (2024) 3D Shape Recognition Multi-view Images Multi-view Images Knowledge distillation with GMViT for efficient 3D shape analysis.
@@
MT-MV-KDF[343] (2024) Myocardial Infarction Detection and Localization Multi-view ECG Single-view ECG Multi-task, multi-view distillation with CNN and attention modules.
V-B1 3D Object Detection
@@
3D object detection has been a focal point in multi-view learning, with numerous studies proposing novel distillation techniques to enhance performance [333, 334, 210, 209, 335, 341, 340]. To this end, SimDistill [341] introduces a LiDAR-camera fusion-based teacher and a simulated fusion-based student for multi-modal learning. By maintaining identical architectures and incorporating a geometry compensation module, the student learns to generate multi-modal features solely from multi-view images.
@@
[210] proposes a comprehensive framework for multi-camera 3D object detection by leveraging cross-task and cross-modal distillation. Cross-task distillation from an instance segmentation teacher avoids ambiguous error propagation, while cross-modal feature distillation and adversarial training refine 3D representations using a LiDAR-based teacher. Cross-modal output distillation further enhances detection accuracy.
@@
[209] proposes a KD framework for aligning a BEV-based student’s features with a LiDAR-based teacher’s, addressing the depth and geometry inference issue compared to LiDAR-based methods.
@@
Cross-modal KD often suffers from feature distribution mismatches. FSD scheme [340] eliminates the need for pre-trained teachers and complex strategies.
@@
BEV-LGKD [333] proposes a LiDAR-guided KD framework for multi-view BEV 3D object detection. By transforming LiDAR points into BEV space and generating foreground masks, the method guides RGB-based BEV models without requiring LiDAR at inference. Depth distillation is also incorporated to improve depth estimation, enhancing BEV perception performance.
@@
BEVDistill [334] introduces a cross-modal BEV distillation framework for multi-view 3D object detection by unifying image and LiDAR features in BEV space and adaptively transfers knowledge across non-homogeneous representations in a teacher-student paradigm.
@@
STXD [335] proposes a structural and temporal cross-modal distillation framework for multi-view 3D object detection. It reduces redundancy in the student’s feature components by regularizing cross-correlation while maximizing cross-modal similarities. Additionally, it encodes temporal relations across frames using similarity maps and employs response distillation to enhance output-level knowledge transfer.
V-B2 3D Shape Recognition
@@
3D shape recognition is another critical area where multi-view distillation has been applied to bridge the gap between 2D and 3D representations [313, 342]. PointMCD [313] bridges 2D visual and 3D geometric domains through a multi-view cross-modal distillation framework, by distilling knowledge from the teacher to a point encoder student. Group Multi-View Transformer (GMViT) [342] addresses the limitations of view-based 3D shape recognition methods, which often struggle with large model sizes that are unsuitable for memory-limited devices. To overcome this, GMViT introduces a large high-performing model designed to enhance the capabilities of smaller student models
V-B3 Other Tasks
@@
In addition to 3D object detection and 3D shape recognition, multi-view distillation techniques have been applied to a range of other tasks, including multi-view image classification [332], vision-based robotic manipulation [336], multi-view BEV detection [337], multi-view stereo depth reconstruction [331], semantic segmentation [338], human action recognition [339], and medical applications [343]. These studies are summarized as follows:
@@
MTS-Net [332] integrates KD with multi-view learning, redefining the roles of the teacher and student models. This end-to-end approach optimizes both view classification and knowledge transfer, with extensions like MTSCNN refining multi-view feature learning for image recognition.
@@
In robotic manipulation, [336] improves vision-based reinforcement learning by transferring knowledge from a teacher policy trained with multiple camera viewpoints to a student policy using a single viewpoint.
@@
[337] uses a spatio-temporal distillation and BEV response distillation by aligning the student’s outputs with those of the teacher, addressing the computational complexity of multi-view BEV detection methods.
@@
Supervised multi-view stereo methods face challenges due to the scarcity of ground-truth depth data. The self-supervised KD-MVS framework [331] uses a teacher trained with photometric and featuremetric consistency to probabilistically transfer knowledge to a student.
@@
MVKD [338] introduces a multi-view KD framework for efficient semantic segmentation. The framework aggregates knowledge from multiple teacher models and transfers it to a student model. To ensure consistency among the multi-view knowledge, MVKD employs a multi-view co-tuning strategy. Additionally, it proposes multi-view feature distillation loss and multi-view output distillation loss to effectively transfer knowledge from multiple teachers to the student.
@@
MKDT [339] introduces a multi-view KD transformer framework for human action recognition, addressing the challenge of incomplete multi-view data. The framework consists of a teacher network and a student network, both utilizing a hierarchical vision transformer with shifted windows to effectively capture spatio-temporal information.
@@
MT-MV-KDF [343] introduces a novel multi-task multi-view KD framework for myocardial infarction detection and localization. Multi-view learning extracts ECG feature representations from different views, while multi-task learning captures similarities and differences across related tasks.
@@
In summary, multi-view distillation techniques have shown significant potential across a broad range of tasks, from 3D object detection and shape recognition to specialized applications such as robotic manipulation and medical diagnostics. The variety of approaches, as highlighted in Table VII, emphasizes the adaptability of these methods to various modalities and challenges.
V-C Text Input
@@
Neural Machine Translation (NMT): While NMT has significantly outperformed traditional statistical methods, large models like transformers are computationally intensive. KD addresses this challenge by distilling the teacher’s output to the student, achieving comparable translation quality with reduced computational costs. This enables the development of more efficient and scalable NMT solutions [344, 345, 346, 347, 348]
@@
Question Answering (QA): In QA, KD transfers both answer predictions and intermediate representations, such as attention distributions, from a large teacher model to a smaller student model. This process helps maintain high accuracy while minimizing computational costs, making QA systems more efficient and suitable for real-world applications [349, 350, 351, 352, 353].
@@
Text Generation: In text generation, KD enables the transfer of a teacher’s generative capabilities to a smaller student model, training it to mimic the vocabulary distributions and language patterns of the teacher. [354] focuses on leveraging BERT’s contextual understanding for efficient text generation, while [355] enhances fluency and diversity by combining distillation with GANs.
@@
Event Detection: KD is applied in event detection to distill knowledge from larger, more complex models to smaller, more efficient ones. [356] improves event detection across languages by distilling knowledge from high-resource to low-resource models, while [357] uses distillation to retain learned knowledge while adapting to new events. Meanwhile, [143] employs adversarial imitation to enhance detection accuracy.
@@
Document Retrieval: In document retrieval, KD transfers the ranking and matching capabilities of large teacher models to smaller student models, typically replicating relevance scores and interaction features [358, 359].
@@
Text Recognition: KD in text recognition facilitates the transfer of both visual and contextual features from a teacher to a student, allowing the student to maintain high recognition accuracy with reduced computational cost [360, 361].
@@
Named Entity Recognition (NER): In NER, KD helps a smaller student model learn from the teacher’s contextual understanding and classification abilities [362, 363].
@@
Text Summarization: KD in text summarization transfers summarization capabilities from teacher to student, training the student to replicate output sequences and probability distributions [364, 365, 366].
@@
Natural Language Understanding (NLU): KD in NLU transfers the rich linguistic understanding of the teacher models to the student. In NLU, the student learns to replicate the teacher’s output probabilities and internal representations across various tasks, ensuring high performance while reducing computational costs [367, 368, 369].
@@
Sentiment Analysis: In sentiment analysis, KD distills the sentiment classification capabilities of the teacher to the student, training it to mimic the teacher’s probability distributions and decision patterns [370, 371].
@@
Text Classification: In text classification, KD transfers classification capabilities from the teacher to the student, enabling the student to mimic the teacher’s feature representations and output probabilities [372, 373].
V-D Speech Input
@@
Speech Recognition (ASR): KD is important in enhancing ASR models by transferring expertise from high-precision teacher models to student models, supporting rapid inference with minimal loss in accuracy. This also improves domain adaptation, making ASR effective in poor acoustic environments or in expert domains, including medical transcription [374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384, 385].
@@
Speech Enhancement: KD in speech enhancement enables denoising through high-fidelity speech representations and model compression [386, 387, 388, 389, 390].
@@
Speaker Recognition and Verification: In speaker recognition and verification, KD helps reduce model complexity while preserving speaker identity features, allowing for fast and secure authentication. It also enhances robustness against adversarial attacks and cross-domain variations, improving performance in noisy or multilingual environments [391, 392, 393, 394, 395, 396, 397].
@@
Speech Translation: KD improves the efficiency of multilingual translation models by enabling knowledge sharing between high-resource and low-resource languages [398, 399, 400, 401]. This enhances translation fluency while maintaining low latency for real-time applications.
@@
Speech Synthesis (Text-to-Speech): In speech synthesis, KD reduces the complexity of deep generative models while retaining natural-sounding speech output [402, 403, 404]. It also helps distill expressive features, making synthetic voices more human-like with lower computation costs.
@@
Speech Separation: KD compresses complex separation models into smaller networks while preserving speaker distinction [405, 406]. It enables real-time speaker and cocktail-party effect reduction on low-power devices.
@@
Spoken Language Identification and Understanding: In this task, KD enables language adaption in low-resource languages through the transfer of information between high-resource multilingual teacher models [407, 408, 409, 410, 411, 412, 413]. As a result, such mechanism empowers lighter, accelerated models to achieve high performance in language identification and speech dialogue comprehension.
@@
Deepfake Speech and Spoofing Detection: With growing concerns about synthetic voice fraud, the need for reliable approaches to deepfake detection has increased. KD strengthens anti-spoofing by transferring knowledge between DNNs trained on a range of deepfake speech datasets and lightweight detection architectures [414, 415, 416, 417, 418, 419].
@@
Audio Classification and Tagging: KD strengthens feature learning relevant to sound event detection, enhancing model efficiency and allowing deployment on embedded platforms such as IoT and edge AI platforms [420, 421, 422, 423, 424, 425].
@@
Spoken Question Answering and Conversational AI: KD enables efficient question-answer through the distillation of contextual information in deep, lengthy transformer models, mapping them onto efficient, quick-answer-producing models [426, 427, 380, 428].
@@
Audio Captioning and Retrieval: In this task, KD enables deep audio-text embedding models to be compressed while maintaining generalization, making them suitable for efficient multimedia search architectures [429, 430].
@@
TABLE VIII: Summary of video knowledge distillation methods based on their sources of distillation.

### v2

@@ 5.2 Multi-view Input
Table 7: Summary of Knowledge Distillation in Multi-View Tasks.
@@
KD-MVS(Ding et al., 2022b) Multi-view Stereo Depth Reconstruction Multi-view Images Multi-view Images Self-supervised MVS distillation for depth reconstruction.
@@
MTS-Net(Tian et al., 2022b) Multi-view Image Classification Multi-view Images Multi-view Images Multi-view learning and distillation for image classification.
@@
BEV-LGKD(Li et al., 2022c) Multi-view BEV 3D Object Detection LiDAR Multi-view RGB Images LiDAR-guided distillation with foreground masks and depth distillation for BEV perception.
@@
BEVDistill(Chen et al., 2022b) Multi-view 3D Object Detection LiDAR Multi-view Images Cross-modal BEV distillation for unifying image and LiDAR features.
@@
STXD(Jang et al., 2023) Multi-view 3D Object Detection LiDAR Multi-view Images Structural and temporal distillation with cross-correlation regularization and response distillation.
@@
PointMCD(Zhang et al., 2023f) 3D Shape Recognition Multi-view Images 3D Point Cloud Multi-view cross-modal distillation for 3D shape recognition.
@@
Acar et al.(Acar et al., 2023) Vision-based Robotic Manipulation Multi-view Images Single-camera Distillation from multi-camera to single-camera for robust manipulation.
@@
Zhang et al.(Zhang et al., 2023e) Multi-view BEV Detection Multi-view BEV Images Multi-view BEV Images Structured distillation for efficient BEV detection.
@@
MVKD(Wang et al., 2023c) Semantic Segmentation Multi-view Images Single-view Images Multi-view distillation with co-tuning and feature/output distillation losses.
@@
MKDT(Lin & Tseng, 2023) Human Action Recognition Multi-view Videos Single-view Videos Multi-view distillation with hierarchical vision transformer for incomplete data.
@@
3X3KD(Klingner et al., 2023) Multi-camera 3D Object Detection LiDAR Multi-view Images Cross-modal distillation for multi-camera 3D object detection.
@@
DistillBEV(Wang et al., 2023p) 3D Object Detection for Autonomous Driving LiDAR Multi-view BEV Images Distillation from LiDAR to BEV for enhanced 3D perception.
@@
FSD-BEV(Jiang et al., 2024e) 3D Object Detection for Autonomous Driving Self-distillation Self-distillation Foreground self-distillation for single-model distillation.
@@
SimDistill(Zhao et al., 2024a) Multi-modal 3D Object Detection LiDAR-camera Fusion Multi-view Images Distillation for 3D object detection with modality gap compensation.
@@
GMViT(Xu et al., 2024b) 3D Shape Recognition Multi-view Images Multi-view Images Knowledge distillation with GMViT for efficient 3D shape analysis.
@@
MT-MV-KDF(Qiang et al., 2024) Myocardial Infarction Detection and Localization Multi-view ECG Single-view ECG Multi-task, multi-view distillation with CNN and attention modules.
5.2.1 3D Object Detection
@@
3D object detection has been a focal point in multi-view learning, with numerous studies proposing novel distillation techniques to enhance performance (Li et al., 2022c; Chen et al., 2022b; Klingner et al., 2023; Wang et al., 2023p; Jang et al., 2023; Zhao et al., 2024a; Jiang et al., 2024e). To this end, SimDistill (Zhao et al., 2024a) introduces a LiDAR-camera fusion-based teacher and a simulated fusion-based student for multi-modal learning. By maintaining identical architectures and incorporating a geometry compensation module, the student learns to generate multi-modal features solely from multi-view images.
@@
Klingner et al. (2023) proposes a comprehensive framework for multi-camera 3D object detection by leveraging cross-task and cross-modal distillation. Cross-task distillation from an instance segmentation teacher avoids ambiguous error propagation, while cross-modal feature distillation and adversarial training refine 3D representations using a LiDAR-based teacher. Cross-modal output distillation further enhances detection accuracy.
@@
Wang et al. (2023p) proposes a KD framework for aligning a BEV-based student’s features with a LiDAR-based teacher’s, addressing the depth and geometry inference issue compared to LiDAR-based methods.
@@
Cross-modal KD often suffers from feature distribution mismatches. FSD scheme (Jiang et al., 2024e) eliminates the need for pre-trained teachers and complex strategies.
@@
BEV-LGKD (Li et al., 2022c) proposes a LiDAR-guided KD framework for multi-view BEV 3D object detection. By transforming LiDAR points into BEV space and generating foreground masks, the method guides RGB-based BEV models without requiring LiDAR at inference. Depth distillation is also incorporated to improve depth estimation, enhancing BEV perception performance.
@@
BEVDistill (Chen et al., 2022b) introduces a cross-modal BEV distillation framework for multi-view 3D object detection by unifying image and LiDAR features in BEV space and adaptively transfers knowledge across non-homogeneous representations in a teacher-student paradigm.
@@
STXD (Jang et al., 2023) proposes a structural and temporal cross-modal distillation framework for multi-view 3D object detection. It reduces redundancy in the student’s feature components by regularizing cross-correlation while maximizing cross-modal similarities. Additionally, it encodes temporal relations across frames using similarity maps and employs response distillation to enhance output-level knowledge transfer.
5.2.2 3D Shape Recognition
@@
3D shape recognition is another critical area where multi-view distillation has been applied to bridge the gap between 2D and 3D representations (Zhang et al., 2023f; Xu et al., 2024b). PointMCD (Zhang et al., 2023f) bridges 2D visual and 3D geometric domains through a multi-view cross-modal distillation framework, by distilling knowledge from the teacher to a point encoder student. Group Multi-View Transformer (GMViT) (Xu et al., 2024b) addresses the limitations of view-based 3D shape recognition methods, which often struggle with large model sizes that are unsuitable for memory-limited devices. To overcome this, GMViT introduces a large high-performing model designed to enhance the capabilities of smaller student models
5.2.3 Other Tasks
@@
In addition to 3D object detection and 3D shape recognition, multi-view distillation techniques have been applied to a range of other tasks, including multi-view image classification (Tian et al., 2022b), vision-based robotic manipulation (Acar et al., 2023), multi-view BEV detection (Zhang et al., 2023e), multi-view stereo depth reconstruction (Ding et al., 2022b), semantic segmentation (Wang et al., 2023c), human action recognition (Lin & Tseng, 2023), and medical applications (Qiang et al., 2024). These studies are summarized as follows:
@@
MTS-Net (Tian et al., 2022b) integrates KD with multi-view learning, redefining the roles of the teacher and student models. This end-to-end approach optimizes both view classification and knowledge transfer, with extensions like MTSCNN refining multi-view feature learning for image recognition.
@@
In robotic manipulation, Acar et al. (2023) improves vision-based reinforcement learning by transferring knowledge from a teacher policy trained with multiple camera viewpoints to a student policy using a single viewpoint.
@@
Zhang et al. (2023e) uses a spatio-temporal distillation and BEV response distillation by aligning the student’s outputs with those of the teacher, addressing the computational complexity of multi-view BEV detection methods.
@@
Supervised multi-view stereo methods face challenges due to the scarcity of ground-truth depth data. The self-supervised KD-MVS framework (Ding et al., 2022b) uses a teacher trained with photometric and featuremetric consistency to probabilistically transfer knowledge to a student.
@@
MVKD (Wang et al., 2023c) introduces a multi-view KD framework for efficient semantic segmentation. The framework aggregates knowledge from multiple teacher models and transfers it to a student model. To ensure consistency among the multi-view knowledge, MVKD employs a multi-view co-tuning strategy. Additionally, it proposes multi-view feature distillation loss and multi-view output distillation loss to effectively transfer knowledge from multiple teachers to the student.
@@
MKDT (Lin & Tseng, 2023) introduces a multi-view KD transformer framework for human action recognition, addressing the challenge of incomplete multi-view data. The framework consists of a teacher network and a student network, both utilizing a hierarchical vision transformer with shifted windows to effectively capture spatio-temporal information.
@@
MT-MV-KDF (Qiang et al., 2024) introduces a novel multi-task multi-view KD framework for myocardial infarction detection and localization. Multi-view learning extracts ECG feature representations from different views, while multi-task learning captures similarities and differences across related tasks.
@@
In summary, multi-view distillation techniques have shown significant potential across a broad range of tasks, from 3D object detection and shape recognition to specialized applications such as robotic manipulation and medical diagnostics. The variety of approaches, as highlighted in Table 7, emphasizes the adaptability of these methods to various modalities and challenges.
5.3 Text Input
@@
Neural Machine Translation (NMT): While NMT has significantly outperformed traditional statistical methods, large models like transformers are computationally intensive. KD addresses this challenge by distilling the teacher’s output to the student, achieving comparable translation quality with reduced computational costs. This enables the development of more efficient and scalable NMT solutions (Kim & Rush, 2016; Tan et al., 2019; Wang et al., 2021a; Liang et al., 2022a; Jin et al., 2024)
@@
Question Answering (QA): In QA, KD transfers both answer predictions and intermediate representations, such as attention distributions, from a large teacher model to a smaller student model. This process helps maintain high accuracy while minimizing computational costs, making QA systems more efficient and suitable for real-world applications (Hu et al., 2018; Izacard & Grave, 2020; Yang et al., 2020; Liu et al., 2021b; Yin et al., 2024a).
@@
Text Generation: In text generation, KD enables the transfer of a teacher’s generative capabilities to a smaller student model, training it to mimic the vocabulary distributions and language patterns of the teacher. Chen et al. (2019c) focuses on leveraging BERT’s contextual understanding for efficient text generation, while Haidar & Rezagholizadeh (2019) enhances fluency and diversity by combining distillation with GANs.
@@
Event Detection: KD is applied in event detection to distill knowledge from larger, more complex models to smaller, more efficient ones. Ren et al. (2021) improves event detection across languages by distilling knowledge from high-resource to low-resource models, while Yu et al. (2021) uses distillation to retain learned knowledge while adapting to new events. Meanwhile, Liu et al. (2019a) employs adversarial imitation to enhance detection accuracy.
@@
Document Retrieval: In document retrieval, KD transfers the ranking and matching capabilities of large teacher models to smaller student models, typically replicating relevance scores and interaction features (Shakeri et al., 2019; Chen et al., 2021c).
@@
Text Recognition: KD in text recognition facilitates the transfer of both visual and contextual features from a teacher to a student, allowing the student to maintain high recognition accuracy with reduced computational cost (Bhunia et al., 2021; Wang & Du, 2021).
@@
Named Entity Recognition (NER): In NER, KD helps a smaller student model learn from the teacher’s contextual understanding and classification abilities (Liang et al., 2021; Ge et al., 2024).
@@
Text Summarization: KD in text summarization transfers summarization capabilities from teacher to student, training the student to replicate output sequences and probability distributions (Chen et al., 2017; Liu et al., 2020b; Nguyen & Luu, 2022).
@@
Natural Language Understanding (NLU): KD in NLU transfers the rich linguistic understanding of the teacher models to the student. In NLU, the student learns to replicate the teacher’s output probabilities and internal representations across various tasks, ensuring high performance while reducing computational costs (Liu et al., 2019c; Tang et al., 2019; FitzGerald et al., 2022).
@@
Sentiment Analysis: In sentiment analysis, KD distills the sentiment classification capabilities of the teacher to the student, training it to mimic the teacher’s probability distributions and decision patterns (Lehečka et al., 2020; Malki, 2023).
@@
Text Classification: In text classification, KD transfers classification capabilities from the teacher to the student, enabling the student to mimic the teacher’s feature representations and output probabilities (Xu & Yang, 2017; Li & Li, 2021).
5.4 Speech Input
@@
Speech Recognition (ASR): KD is important in enhancing ASR models by transferring expertise from high-precision teacher models to student models, supporting rapid inference with minimal loss in accuracy. This also improves domain adaptation, making ASR effective in poor acoustic environments or in expert domains, including medical transcription (Chebotar & Waters, 2016; Markov & Matsui, 2016; Takashima et al., 2018; Mun’im et al., 2019; Kurata & Saon, 2020; Yoon et al., 2021; You et al., 2021a; Yoon et al., 2022; Zhao et al., 2022c; Kim et al., 2023b; Park et al., 2023; Zhang et al., 2023j).
@@
Speech Enhancement: KD in speech enhancement enables denoising through high-fidelity speech representations and model compression (Watanabe et al., 2017; Wu et al., 2019; Kim & Kim, 2021; Shin et al., 2023; Park et al., 2024).
@@
Speaker Recognition and Verification: In speaker recognition and verification, KD helps reduce model complexity while preserving speaker identity features, allowing for fast and secure authentication. It also enhances robustness against adversarial attacks and cross-domain variations, improving performance in noisy or multilingual environments (Mingote et al., 2020; Peng et al., 2022; Liu et al., 2022a; Jin et al., 2023b; Mingote et al., 2023; Truong et al., 2024; Hoang et al., 2024).
@@
Speech Translation: KD improves the efficiency of multilingual translation models by enabling knowledge sharing between high-resource and low-resource languages (Liu et al., 2019e; Gaido et al., 2020; Inaguma et al., 2021; Lei et al., 2023). This enhances translation fluency while maintaining low latency for real-time applications.
@@
Speech Synthesis (Text-to-Speech): In speech synthesis, KD reduces the complexity of deep generative models while retaining natural-sounding speech output (Oord et al., 2018; Xu et al., 2020a; Li et al., 2024f). It also helps distill expressive features, making synthetic voices more human-like with lower computation costs.
@@
Speech Separation: KD compresses complex separation models into smaller networks while preserving speaker distinction (Chen et al., 2018; Zhang et al., 2021b). It enables real-time speaker and cocktail-party effect reduction on low-power devices.
@@
Spoken Language Identification and Understanding: In this task, KD enables language adaption in low-resource languages through the transfer of information between high-resource multilingual teacher models (Shen et al., 2018; 2019c; 2020; Kim et al., 2021b; Cappellazzo et al., 2022; Mao & Zhang, 2023; Cappellazzo et al., 2023). As a result, such mechanism empowers lighter, accelerated models to achieve high performance in language identification and speech dialogue comprehension.
@@
Deepfake Speech and Spoofing Detection: With growing concerns about synthetic voice fraud, the need for reliable approaches to deepfake detection has increased. KD strengthens anti-spoofing by transferring knowledge between DNNs trained on a range of deepfake speech datasets and lightweight detection architectures (Liao et al., 2022; Xue et al., 2023; Ren et al., 2023c; b; Lu et al., 2024; Wani & Amerini, 2025).
@@
Audio Classification and Tagging: KD strengthens feature learning relevant to sound event detection, enhancing model efficiency and allowing deployment on embedded platforms such as IoT and edge AI platforms (Chang et al., 2020; Yin et al., 2021; Schmid et al., 2023; Liang et al., 2022b; Dinkel et al., 2023; Tang et al., 2024).
@@
Spoken Question Answering and Conversational AI: KD enables efficient question-answer through the distillation of contextual information in deep, lengthy transformer models, mapping them onto efficient, quick-answer-producing models (You et al., 2020a; b; 2021a; 2021b).
@@
Audio Captioning and Retrieval: In this task, KD enables deep audio-text embedding models to be compressed while maintaining generalization, making them suitable for efficient multimedia search architectures (Xu et al., 2024c; Primus & Widmer, 2024).
@@
Table 8: Summary of video knowledge distillation methods based on their sources of distillation.